The results of the study are presented in the table below.

Significant Reduction in Pain? Type of Injection NO YES Total SALINE 19 3 22 BOTOX 11 10 21 Total 30 13 43

7. What is the probability a patient in this experiment had a significant reduction in pain? Write the fraction and the decimal. Round the decimal to two places. P(YES) =

8. What is the probability a patient has a Botox injection AND had a significant reduction in pain? Write the fraction and the decimal. Round the decimal to two places. P(BOTOX and YES) =

9. What is the probability that a randomly selected patient that received a saline injection had a significant reduction in pain? Write the fraction and the decimal. Round the decimal to two places. P(YES | SALINE) =

10. What is the probability that a patient who received a Botox injection had a significant reduction in pain? Write the fraction and the decimal. Round the decimal to two places. P(YES | BOTOX) =

11. Are the type of injection and having a significant reduction in pain independent or dependent? Explain your answer.

12. Are having a saline injection and having a significant reduction in pain mutually exclusive?
Answer #1

7) Probability that a patient in this experiment had a significant reduction in pain , P(YES) = 13/43 = 0.30

8) P(BOTOX and YES) = 10/43 = 0.23

9) P(YES | SALINE) = P(YES and SALINE)/P(SALINE) = 3/22 = 0.14

10) P(YES | BOTOX) = P(YES and BOTOX)/P(BOTOX)

= 10/21 = 0.48

11) P(significant reduction in pain and type of injection as SALINE)

= P(YES and SALINE) = 3/43 = 0.07

P(YES)*P(SALINE) = 13/43 * (22/43) = 0.15

Since, these are not equal .The type of injection and having a significant reduction in pain is dependent.

12) Since, P(SALINE and YES) = 0.07 which is not equal to 0, thus these are not mutually exclusive.